# Quantum Monte Carlo description of correlated electrons in two-dimensional FeSe

### v1

@@ Abstract
Electronic correlation effects on the structural properties of double-layer FeSe are studied using variation and diffusion quantum Monte Carlo methods. The Slater-Jastrow many-body wavefunction with two different forms for the homogeneous two-body pair-correlation term is used. The ground-state energy of the system is obtained at the thermodynamic limit using two different trial wave functions called JDFT and JSD. Only the Jastrow factor is fully optimized in the JDFT wave function, while the Slater determinant comes from the density functional approximation. In the JSD trial wave function, the Slater determinant and the Jastrow factor are fully optimized simultaneously. We calculated the VMC and DMC energies as a function of interlayer separation for two different in-plane iron-iron bond lengths. Our QMC results indicate that the optimized interlayer separation decreases with increasing iron-iron bond length driven by stretch. We found that a three- to two-dimensional phase transition increases the electron-electron correlation effects and shifts the system from moderately correlated to strongly correlated. The value of correlation energy implies that the Hubbard $U$ value, which is widely used in DFT+U calculations, for two-dimensional FeSe is larger than its value in bulk.
@@ I Introduction
The electron-electron (el-el) many-body correlations in magnetic two-dimensional (2D) systems, including iron-based superconductors (FeSCs), play a fundamental role in understanding their structural, magnetic and exotic electronic properties such as unconventional superconductivity [1, 2, 3, 4, 5, 6, 7]. In two-dimensional FeSCs, electronic interactions are moderately strong, placing them in an intermediate coupling regime between weak coupling and strong correlation effects found in cuprates[8]. The confinement of electrons in two-dimensional FeSCs enhances correlation effects due to the reduced screening of Coulomb interactions between electrons that introduce magnetic anisotropy, which can be tuned externally by electric fields, strain, photoexcitation, and chemical doping[9, 10, 11, 12, 4, 13]. Renormalization of the electronic band structure and the increase in the effective mass of electrons, which is observed in angle-resolved photoemission spectroscopy (ARPES) and quantum oscillation experiments[14], can be enhanced by the el-el correlation interaction. Deviations from Fermi-liquid behavior [15, 16] in FeSCs are introduced by electronic correlations, especially near quantum critical points where magnetic fluctuations dominate. This non-Fermi liquid behavior is reflected in unusual temperature-dependent properties $\rho\sim T^{n},n>2$ and other anomalous properties [17]. Competing orders, such as charge density waves or nematic order, which often coexist or compete with superconductivity, can be stabilized due to correlation effects[18, 19, 20]. Nematicity, characterized by the break of rotational symmetry, is strongly influenced by many-body el-el correlations in FeSCs[21]. Two-dimensional FeSCs provide an ideal platform for studying the interplay between electron correlations and reduced dimensionality, bridging the gap between weakly correlated systems (such as conventional superconductors) and strongly correlated systems (such as cuprates) [10, 2, 1, 4]. In this work, many-body el-el correlations in two-dimensional iron selenide (2D-FeSe) are calculated using continuous quantum Monte Carlo methods in real space.
@@
In this work, we focus on double-layer 2D-FeSe, its electronic properties can be influenced by interlayer separation, which is governed by long-range van der Waals (vdW) interactions [33]. Controlling this separation, for example, through isothermal compression, allows tuning of various emergent features, such as the coupling between nematicity and magnetism [12]. The coexistence of competing orders, namely weak vdW forces and strong localized interactions in double-layer 2D-FeSe system, presents unique opportunities to explore exotic quantum phenomena, with implications for both fundamental science and technological applications. Our main tools are real space variation and diffusion quantum Monte Carlo methods which can accurately describe the correlation driven phenomena.
Real-space Variational Quantum Monte Carlo (VMC) and Diffusion Quantum Monte Carlo (DMC) have proven to be highly successful in studying electronic correlation effects in transition metal compounds and capturing weak vdW interaction[34, 35, 36, 37], including systems like FeSe[38], due to the fact that they directly address the many-body nature of electron interactions, which are critical in correlated systems with partially filled $d$-orbitals. Traditional mean-field approaches, such as density functional theory (DFT) with local density approximation (LDA) or generalized gradient approximation (GGA), often fail to capture strong correlations. VMC and DMC go beyond mean-field approximations by directly solving the many-body Schrödinger equation using the stochastic techniques. Both VMC and DMC allow for systematic improvement of accuracy by refining trial wave functions and optimizing variational parameters using variance and energy minimization[39].
@@ II Method
The main ingredient of our QMC calculations is the many-body wavefunction with the resonance value bond (RVB) format defined as the product of a Jastrow factor $J$ and an antisymmetrized geminal power (AGP) determinant part $\Psi_{\text{AGP}}$ [43, 27] as implemented in TurboRVB[44]. The determinant part is:
@@
$\phi(\mathbf{r}^{\uparrow},\mathbf{r}^{\downarrow})=\sum_{i=1}^{N/2}\alpha_{i}\psi_{i}^{MO}(\mathbf{r}^{\uparrow})\psi_{i}^{MO}(\mathbf{r}^{\downarrow}),$ (2)
The MOs are expanded in a Gaussian single-particle basis set ${\chi}$ centered on the atomic position $\psi_{i}^{MO}(\mathbf{r})=\sum_{j}\beta_{ij}\chi_{j}(\mathbf{r})$ [43, 45]. If not mentioned, our results are obtained using an uncontracted Gaussian basis of $8s6p4d1f$ and $6s4p2d$ orbitals for Fe and Se, respectively. Initial values of the Gaussian orbitals chosen from the cc-pVTZ [46] basis set. The core electrons of the Fe and Se atoms were replaced by correlation consistent effective core potentials (ccECPs)[47, 48]. Our simulation cell is subject to two-dimensional periodic boundary conditions in the $xy$ direction with standard Ewald summations to calculate the Coulomb interaction. The variational parameters in our $AGP$ wave function are $\alpha_{i}$, $\beta_{ij}$, and the exponents of the uncontracted Gaussian basis set ${\chi}$. The initial MOs were obtained using the density functional approach [49] with the local density approximation (LDA) [50] using the same uncontracted Gaussian basis set described above.
@@
where $g_{lm}^{ab}$ are optimizable parameters. The three-body electron-ion-electron is defined by the diagonal matrix elements $g^{aa}$. The off-diagonal $g^{ab},a\neq b$ matrix elements define the four-body electron-ion-electron-ion terms that were not used in our study. $f(\mathbf{r}_{i},\mathbf{r}_{j})$ is a two-electron coordinate function expanded using one-particle basis sets, $i,j$ are electron indices, and $r_{ij}$ are electron-electron distances. We used two different forms for the two-body homogeneous part $u_{F}=r/(2(1+ar))$ [51] and $u_{C}=\frac{b}{2}(1-exp(-r/b))$ [52], where $r$ is the relative distance between two electrons, $a$ and $b$ are variational parameters. For Jastrow single-particle orbitals, we used uncontracted Gaussian basis sets of $3s3p1d$ and $3s2p$ for Fe and Se, respectively. The DMC energies are obtained using the time step 0.01 a.u., the 3200 number of walkers (configurations) and the locality approximation for the pseudopotential[53].
III Numerical Results
Figure 1 shows the difference between the VMC and DFT energies during the last 100 optimization steps out of several hundreds of optimization iterations. The number of optimization steps was increased by system size. The results of the WF optimization procedure are shown for three system sizes and two-body Jastrow forms of $u_{F}$ and $u_{C}$, which are defined in Section II. Only the coefficients of Jastrow factor were optimized in Fig. 1(a), while Fig. 1(b) shows the results for full optimization of the Jastrow term. The Jastrow term and the Slater determinant were optimized simultaneously in Fig. 1(c). The VMC and DMC energies obtained by the wave functions $JDFT$ and $JSD$ as a function of the system size are shown in Fig. 2. All the results of Figs. 1 and 2 were obtained for double layer FeSe with in-plane iron-iron bond length 3.689 Åand interlayer separation 5.83 Å, which is the optimized value predicted by our DMC results (Fig. 3). The ground state energy of the system at the infinite system size limit is obtained by using linear extrapolation of the VMC and DMC energies as a function of $1/N$, where $N$ is the number of electrons in the simulation cell. The values of VMC and DMC energies at each system size and the thermodynamic limit are presented in Table 1.
@@ III Numerical Results
Figure 1: The difference between VMC and DFT energies as a function of the number of optimization interation. (a) Only the coefficients of the Jastrow terms are optimised, (b) The exponent and Coefficients of the Jastrow term are optimized simultaneously, (c) The Jastrow and Slater determinat are optimised. Two Jastrow forms of $u_{F}$ and $u_{C}$ were used. Only the last one hundred optimization steps are shown. Three system sizes with number of electrons in simulation cell $N_{el}=176,396,704$ were considered.
$N$ $VMC$-$J_{F}DFT$ $DMC$-$J_{F}DFT$ $VMC$-$J_{F}SD$ $DMC$-$J_{F}SD$ $VMC$-$J_{C}DFT$ $DMC$-$J_{C}DFT$ $VMC$-$J_{C}SD$ $DMC$-$J_{C}SD$
$176$ $-6.022123(8)$ $-6.03398(3)$ $-6.029479(6)$ $-6.03741(3)$ $-6.021325(8)$ $-6.033809(4)$ $-6.029264(6)$ $-6.03740(2)$
$396$ $-6.020526(9)$ $-6.03299(5)$ $-6.027806(7)$ $-6.03603(4)$ $-6.019759(9)$ $-6.032964(6)$ $-6.027308(7)$ $-6.03624(4)$
$704$ $-6.02007(1)$ $-6.03254(8)$ $-6.026830(9)$ $-6.03519(8)$ $-6.01904(2)$ $-6.03235(8)$ $-6.02627(1)$ $-6.03547(9)$
$\infty$ $-6.01934(1)$ $-6.0321(1)$ $-6.02611(1)$ $-6.0346(2)$ $-6.01834(3)$ $-6.0320(2)$ $-6.0232(1)$ $-6.0349(3)$
@@
Table 1: VMC and DMC energies at different system size and the thermodynamic limit obtained using $J_{F}$ and $J_{C}$ two-body el-el Jastrow term. The energy at the infinite system size limit is obtained using the linear extrapolation (Fig. 2).
@@
Figure 2: VMC $E_{VMC}$ (left panel) and DMC $E_{DMC}$ (right panel) energy of double layer FeSe as a function of $1/N$, where $N$ is the number of electrons in the simulation cell, obtained using $J_{F}$ and $J_{C}$ two-body el-el Jastrow factors.
The nearly degenerate electronic configurations close to the minimum of VMC and DMC energy curves (Fig. 3) suggest strong static correlation and multi-reference character of the system driven by strong el-el interactions. This implies the importance of using the multideterminant trial wave function to accurately describe the ground-state electronic structure. To investigate this behavior further, we increased the number of Fe-$d$ orbitals in the single-particle basis set of the Slater determinant and the Jastrow factor. The VMC and DMC energies of the double layer FeSe, which are calculated using the determinant and Jastrow base set of $8s6pnd1f,(n=4,6)$ and $3s3pnd,(n=1,3)$, respectively, for the iron atom, are listed in table 2. The energies are obtained for double-layer FeSe with Fe-Fe bond length of 3.689 Åand interlayer separation of 5.83 Å. The dynamic correlation captured by the Jastrow factor is weaker than the static correlation. Hence, increasing the number of Fe-$d$ orbitals in the slater determinant basis set improves the ground-state energy more than the Jastrow basis set. Compared with DMC, the VMC energies are more affected by the number of Fe-$d$ orbitals in the base set. More importantly, full optimization of the wave function reduces the effect of number of Fe-$d$ orbitals in the basis set as shown by the $DMC$-$JSD$ energies (tab. 2). Comparing the $DMC$-$JSD$ energies of the smallest and largest basis sets shows that reaching the mHa/el accuracy is possible by full optimization of the wave function and without increasing the size of the basis set.

### v2

@@ Abstract
An interesting question in physics is how the correlation energy of atoms evolves upon forming a solid. Here, we address this problem for a specific case of double-layer FeSe. We used many-body wavefunction-based quantum Monte Carlo (QMC) techniques to compute the correlation energies of double-layer FeSe with different geometrical configurations and compared them with those of isolated Fe and Se atoms. Variational and diffusion QMC calculations were carried out with Slater–Jastrow trial wavefunctions, employing two alternative forms for the homogeneous two-body pair-correlation term. The ground-state energy was obtained in the thermodynamic limit using two types of trial wave functions of JDFT, in which only the Jastrow factor is optimized while the Slater determinant is derived from the local density approximation, and JSD, where both the Jastrow factor and the Slater determinant are optimized simultaneously. Our results indicate that the correlation energy of double layer FeSe at the thermodynamic limit is mainly determined by the atomic contributions, with the bonding between atoms playing a comparatively minor role in it. After optimizing the interlayer separation of double-layer FeSe under tensile strain, we analyze the correlation energy as a function of strain and separation. We found that with increasing tensile stretch and interlayer spacing, the correlation energy of double-layer FeSe stochastically approaches that of its constituent atomic fragments.
@@ I Introduction
The electron-electron (el-el) many-body correlations in magnetic two-dimensional (2D) systems, including iron-based superconductors (FeSCs), play a fundamental role in understanding their structural, magnetic and exotic electronic properties such as unconventional superconductivity [1, 2, 3, 4, 5, 6, 7]. In two-dimensional FeSCs, electronic interactions are moderately strong, placing them in an intermediate coupling regime between weak coupling and strong correlation effects found in cuprates[8]. The confinement of electrons in two-dimensional FeSCs enhances correlation effects due to the reduced screening of Coulomb interactions between electrons that introduce magnetic anisotropy, which can be tuned externally by electric fields, strain, photoexcitation, and chemical doping[9, 10, 11, 12, 4, 13]. Renormalization of the electronic band structure and the increase in the effective mass of electrons, which is observed in angle-resolved photoemission spectroscopy (ARPES) and quantum oscillation experiments[14], can be enhanced by the el-el correlation interaction. Deviations from Fermi-liquid behavior [15] in FeSCs are introduced by electronic correlations, especially near quantum critical points where magnetic fluctuations dominate. This non-Fermi liquid behavior is reflected in unusual temperature-dependent properties $\rho\sim T^{n},n>2$ and other anomalous properties [16]. Competing orders, such as charge density waves or nematic order, which often coexist or compete with superconductivity, can be stabilized due to correlation effects[17, 18, 19]. Nematicity, characterized by the break of rotational symmetry, is strongly influenced by many-body el-el correlations in FeSCs[20]. Two-dimensional FeSCs provide an ideal platform for studying the interplay between electron correlations and reduced dimensionality, bridging the gap between weakly correlated systems (such as conventional superconductors) and strongly correlated systems (such as cuprates) [10, 2, 1, 4]. In this work, many-body el-el correlations in two-dimensional iron selenide (2D-FeSe) are calculated using continuous quantum Monte Carlo methods in real space.
@@
In this work, we focus on double-layer 2D-FeSe. The electronic properties of double-layer FeSe can be influenced by interlayer separation, which is governed by long-range van der Waals (vdW) interactions [32]. Controlling this separation, for example, through isothermal compression, allows tuning of various emergent features, such as the coupling between nematicity and magnetism [12]. The coexistence of competing orders, namely strong localized interactions and weak vdW forces in the double-layer FeSe system, presents unique opportunities to explore exotic quantum phenomena, with implications for both fundamental science and technological applications. Tensile strain applied within the Fe-Se plane can affect the electronic structure and electron correlation in double-layer FeSe. This tuning mechanism alters orbital overlap, bandwidth, and hybridization, making it a useful tool for manipulating correlated phases in double-layer FeSe. Our main tools in this study are the real space variational and diffusion quantum Monte Carlo methods, which can accurately describe the correlation-driven phenomena.
Real-space Variational Quantum Monte Carlo (VMC) and Diffusion Quantum Monte Carlo (DMC) have proven to be highly successful in studying electronic correlation effects in transition metal compounds and capturing weak vdW interaction in 2D systems [33, 34, 35, 36, 37], including systems like FeSe[38], due to the fact that they directly address the many-body nature of electron interactions, which are critical in correlated systems with partially filled $d$ orbitals. Traditional mean-field approaches, such as density functional theory (DFT) with local density approximation (LDA) or generalized gradient approximation (GGA), often fail to capture strong correlations. VMC and DMC go beyond mean-field approximations by directly solving the many-body Schrödinger equation using the stochastic techniques. Both VMC and DMC allow for systematic improvement of accuracy by refining trial wave functions and optimizing variational parameters using variance and energy minimization[39].
@@
We present the results which are obtained using both the VMC and DMC methods. VMC provides a controlled variational estimate of the ground-state energy using trial wave functions with systematically tunable accuracy. It allows us to explore how different forms and parameterizations of the wave function affect observables, making it ideal for analyzing the sensitivity of the correlation energy to nodal structure, electron localization, and bond-length variation. DMC, by projecting out the ground state (within the fixed-node approximation), provides more accurate estimates of total and correlation energies, less sensitive to the details of the trial wave function beyond the nodal surface. It serves as our primary source of benchmark-quality results. Using both methods allows us to (i) validate the robustness of physical trends across two levels of theory, (ii) quantify the impact of wave function quality on the results, and (iii) assess the affect of the nodal surface on QMC results.
@@ II Method
We provide details of our QMC simulations for double-layer FeSe in this section. The main ingredient of our QMC calculations is the many-body wavefunction with the resonance valence bond (RVB) format defined as the product of a Jastrow factor $J$ and an antisymmetrized geminal power (AGP) determinant part $\Psi_{\text{AGP}}$ [43, 26] as implemented in TurboRVB[44]. The determinant part is:
@@
$\phi(\mathbf{r}^{\uparrow},\mathbf{r}^{\downarrow})=\sum_{i=1}^{N/2}\psi_{i}^{MO}(\mathbf{r}^{\uparrow})\psi_{i}^{MO}(\mathbf{r}^{\downarrow}),$ (2)
The MOs are expanded in a Gaussian single-particle basis set $\chi$} centered on the atomic position $\psi_{i}^{MO}(\mathbf{r})=\sum_{j}\beta_{ij}\chi_{j}(\mathbf{r})$ [43, 45]. If not mentioned, our results are obtained using an uncontracted Gaussian basis of $8s6p4d1f$ and $6s4p2d$ orbitals for Fe and Se, respectively. Initial values of the Gaussian orbitals chosen from the cc-pVTZ [46] basis set. The core electrons of the Fe and Se atoms were replaced by correlation consistent effective core potentials (ccECPs)[47, 48]. Our simulation cell is subject to two-dimensional periodic boundary conditions in the $xy$ direction with standard Ewald summations and the $\Gamma$-point to calculate the Coulomb interaction. The variational parameters in our $AGP$ wave function, which is in fact the single Slater determinant, as the maximum number of MOs equals the number of electron pairs, are $\beta_{ij}$, and the exponents of the uncontracted Gaussian basis set $\chi$}. The initial MOs were obtained using the density functional approach [49] with the local density approximation (LDA) [50] using the same uncontracted Gaussian basis set described above.
@@
where $g_{lm}^{ab}$ are optimizable parameters. The three-body electron-ion-electron is defined by the diagonal matrix elements $g^{aa}$. The off-diagonal $g^{ab},a\neq b$ matrix elements define the four-body electron-ion-electron-ion terms that were not used in our study. $f(\mathbf{r}_{i},\mathbf{r}_{j})$ is a two-electron coordinate function expanded using one-particle basis sets, $i,j$ are electron indices, and $r_{ij}$ are electron-electron distances. For FeSe simulations, we used two different forms for the two-body homogeneous part $u_{F}=r/(2(1+ar))$ [51] and $u_{C}=\frac{b}{2}(1-exp(-r/b))$ [52], where $r$ is the relative distance between two electrons, $a$ and $b$ are variational parameters. For Jastrow single-particle orbitals, we used uncontracted Gaussian basis sets of $3s3p1d$ and $3s2p$ for Fe and Se, respectively. The DMC energies are obtained using the time step 0.01 a.u., the 3200 walkers (configurations), and the locality approximation for the pseudopotential[53].
III Results and discussions
III.1 Correlation energy of Fe and Se atoms
To validate our HF approach, we calculated the correlation energy $E_{corr}=E_{QMC}-E_{HF}$ of the Fe and Se pseudo-atoms, where $E_{QMC}$ is the total energy of the atom obtained by the VMC and DMC methods, and $E_{HF}$ is the HF energy calculated by solving the Kohn-Sham Hamiltonian with %100 exchange interaction and %0 correlation weight factor [49]. We used large Gaussian basis sets of $10s10p7d2f1g$ and $9s9p6d1f$ for Fe ans Se atoms, respectively, to reach the complete base set (CBS) limit. We performed QMC calculations using the same basis set as the CBS-HF and also the basis set used for the FeSe calculations ($8s6p4d1f$ and $6s4p2d$ for Fe and Se, respectively) reported in the next section. The Slater determinant was obtained using both LDA and HF methods and $u_{F}$ was used for the two-body homogeneous part of the Jastrow term.
Our results for the Fe atom are listed in Tab. 1. Only the Jastrow term was optimized in the JDFT and JHF WFs. The results in Tab. 1 were obtained using uncontracted Gaussian basis sets of $3s3p1d$ for the Jastrow factor which is the same as the FeSe system.
BS-WF VMC-JDFT VMC-JSD DMC-JDFT DMC-JSD
$8s6p4d1f$-LDA -50.94(5) -62.38(4) -65.04(6) -73.17(6)
$10s10p7d2f1g$-LDA -67.95(1) -68.07(1) -72.42(2) -72.69(2)
BS-WF VMC-JHF VMC-JSD DMC-JHF DMC-JSD
$8s6p4d1f$-HF -50.02(4) -66.14(3) -64.92(6) -74.72(6)
$10s10p7d2f1g$-HF -67.37(2) -67.80(2) -72.42(3) -72.69(3)
Table 1: The correlation energy of Fe pseudoatom (ccECP) in mHa/el obtained using $8s6p4d1f$ and $10s10p7d2f1g$ (CBS) basis sets, and LDA/HF generated Slater determinant for the WF. The correlation energy of all electron atom obtained by RPA and QC methods are -68.2 [54], and -47.5 [55], respectively. The DMC energies were calculated using time step $0.01$ a.u.
Our QMC results for the correlation energy of the Fe pseudoatom agree relatively well with the correlation energy of the all-electron Fe atom obtained by the Random Phase Approximation (RPA) and Quantum Chemistry (QC) methods: -68.2 [54] and -47.5 [55], respectively. The DMC-JSD-$E_{corr}$ results show a negligible dependence on the size of the basis set and whether LDA or HF is used to generate the Slater determinant. We investigated the DMC time-step error using $8s6p4d1f$-HF WF. Fig.1 shows that the DMC time-step error in the $E_{corr}$ results reported in Tab. 1 is negligible. The DMC correlation energies at the zero time step obtained by the JHF and JSD WFs are -64.0(1) and -74.14(6) mHa/el, respectively. We used the same DMC time step of $0.01$ a.u. for the following FeSe calculations.
Figure 1: DMC correlation energy of Fe as a function of time step. Energies were calculated using $8s6p4d1f$-HF WF and ccECP.
The other factor which may affect the QMC correlation energy is PP. To investigate this factor, we calculated the VMC and DMC correlation energy of the Fe atom using $8s6p4d1f$-HF WF and different PPs including ccECP[47, 48], CEPP[56], eCEPP[57], and ECP-Burkatzki[58, 59]. The results, which are listed in Tab. 2, indicate negligible differences between the QMC correlation energies obtained using the studied PPs.
$P$ VMC-JHF VMC-JSD DMC-JHF DMC-JSD
ccECP -50.02(4) -66.14(3) -64.92(6) -74.72(6)
CEPP -52.13(3) -65.65(2) -64.60(6) -72.71(4)
eCEPP -51.33(3) -65.15(2) -64.08(6) -72.56(4)
ECP -54.98(3) -66.32(2) -67.03(6) -74.19(4)
Table 2: Correlation energy of Fe pseudoatom in mHa/el obtained using VMC and DMC with different PPs and $8s6p4d1f$-HF WF.
The VMC and DMC correlation energies of the Se atom using ccECP and LDA-generated WF with two different basis sets are listed in Tab. 3. The correlation energies of the Se all-electron atom obtained by the RPA and QC methods are -71.5 [54] and -51.2 [55], respectively. The VMC/DMC-JSD results are close to the QC all electron results.
BS-WF VMC-JDFT VMC-JSD DMC-JDFT DMC-JSD
$6s4p2d$-LDA -14.34(3) -44.78(1) -15.07(6) -44.76(3)
$9s9p6d1f$-LDA -44.75(2) -45.05(1) -44.78(3) -45.04(2)
Table 3: The correlation energy of Se pseudoatom (ccECP) in mHa/el. The correlation energy of all electron atom obtained by RPA and QC methods are -71.5 [54], and -51.2 [55], respectively. The DMC energies were calculated using time step $0.01$ a.u.
As we mentioned above, all the results $E_{corr}$ for the Fe and Se atoms were calculated using a large basis set to obtain the HF energy at CBS limit. In principle, a larger basis set can be used, which would affect $E_{corr}$. Due to technical difficulties, we only used Gaussian exponents smaller than 100 for CBS-HF calculations. The rest of this paper uses ccECP and the same CBS for HF calculations. The total VMC, DMC, and HF energies of Fe and Se atoms obtained using JDFT and JSD wavefunctions with different PPs are reported in the supplementary material[60].
III.2 Correlation energy of Double-layer FeSe at the thermodynamic limit
In this section, the correlation energy of double-layer FeSe is evaluated in the thermodynamic limit by computing the ground-state total energies of three finite systems using QMC and HF, followed by extrapolation to the infinite-size limit.
We focused on optimizing the trial wave function for each system size and lattice parameter before performing the final VMC and DMC simulations. The trial wave function is crucial in QMC simulations because it directly affects the accuracy, efficiency, and stability of the simulations, as it serves as a guiding function for importance sampling in DMC, which modifies the diffusion equation to improve convergence. A poorly optimized trial wave function can lead to large statistical fluctuations, making it difficult to obtain precise results. Optimizing the trial wave function reduces variance, leading to faster and more stable simulations. We systematically optimized the trial wave function by following three steps: (I) The Jastrow coefficients and the two-body term $u$ were optimized, as shown in Fig. 2(a). In this step, depending on the size of the system, $\sim$ 4-8 mHa/el VMC energy obtained with respect to DFT. (II) The Jastrow function was fully optimized (Fig. 2(b)) and the wave function was called JDFT. The energy difference between VMC and DFT is $\sim$ 8-11 mHa/el, depending on the system size. (III) The Jastrow terms and the Slater determinant were optimized simultaneously, and the wave function was named JSD. Since the FS errors normally increase by the number of electrons in the simulation cell, the VMC energy gain with respect to DFT is smaller for larger systems. We found that following these three steps in order, meaning that the well-optimized wave function at the end of each step is used as the initial wave function for the next step, is required to reach an accuracy of mHa/el and avoid local minima.
Refer to caption Refer to caption Refer to caption
@@ III.2 Correlation energy of Double-layer FeSe at the thermodynamic limit
Figure 2: The difference between VMC and DFT energies as a function of the number of optimization interation. (a) Only the coefficients of the Jastrow terms are optimized, (b) The exponent and Coefficients of the Jastrow term are optimized simultaneously, (c) The Jastrow and Slater determinat are optimised. Two Jastrow forms of $u_{F}$ and $u_{C}$ were used. Only the last one hundred optimization steps are shown. Three system sizes with the number of electrons in the simulation cell $N_{el}=176,396,704$ were considered.
Figure 2 shows the difference between the VMC and DFT energies during the last 100 optimization steps out of several hundreds of optimization iterations. The number of optimization steps was increased by system size. The results of the WF optimization procedure are shown for three system sizes and two-body Jastrow forms of $u_{F}$ and $u_{C}$, which are defined in Section II. The VMC and DMC energies obtained by the wave functions $JDFT$ and $JSD$ as a function of the system size are shown in Fig. 3. All the results of Figs. 2 and 3 were obtained for double layer FeSe with in-plane iron-iron bond length 3.69 Åand interlayer separation 5.83 Å , which is the optimized value predicted by our DMC results (Fig. 4). The ground state energy of the system at the infinite system size limit is obtained by using linear extrapolation of the VMC and DMC energies as a function of $1/N$, where $N$ is the number of electrons in the simulation cell. The values of the VMC and DMC energies at each system size and the thermodynamic limit are presented in Table 4.
NN VMC-$J_{F}DFT$ DMC-$J_{F}DFT$ VMC-$J_{F}SD$ DMC-$J_{F}SD$ VMC-$J_{C}DFT$ DMC-$J_{C}DFT$ VMC-$J_{C}SD$ DMC-$J_{C}SD$ HF-CBS
$176$ 6.022123(8)-6.022123(8) 6.03398(3)-6.03398(3) 6.029479(6)-6.029479(6) 6.03741(3)-6.03741(3) 6.021325(8)-6.021325(8) 6.033809(4)-6.033809(4) 6.029264(6)-6.029264(6) 6.03740(2)-6.03740(2) 5.970143-5.970143
$396$ 6.020526(9)-6.020526(9) 6.03299(5)-6.03299(5) 6.027806(7)-6.027806(7) 6.03603(4)-6.03603(4) 6.019759(9)-6.019759(9) 6.032964(6)-6.032964(6) 6.027308(7)-6.027308(7) 6.03624(4)-6.03624(4) 5.971588-5.971588
$704$ 6.02007(1)-6.02007(1) 6.03254(8)-6.03254(8) 6.026830(9)-6.026830(9) 6.03519(8)-6.03519(8) 6.01904(2)-6.01904(2) 6.03235(8)-6.03235(8) 6.02627(1)-6.02627(1) 6.03547(9)-6.03547(9) 5.973026-5.973026
$\infty$ 6.01934(1)-6.01934(1) 6.0321(1)-6.0321(1) 6.02611(1)-6.02611(1) 6.0346(2)-6.0346(2) 6.01834(3)-6.01834(3) 6.0320(2)-6.0320(2) 6.0232(1)-6.0232(1) 6.0349(3)-6.0349(3) 5.9736-5.9736
@@
Table 4: VMC and DMC energies at different system size and the thermodynamic limit obtained using $J_{F}$ and $J_{C}$ two-body el-el Jastrow term. The energy at the infinite system size limit is obtained using the linear extrapolation (Fig. 3). The last column lists HF energy with CBS.
Performing QMC simulations in a finite supercell with periodic boundary conditions (PBC) suffers from finite-size (FS) errors [61]. The main source of FS errors includes (i) the single-particle FS effect, which is because of the discrete $k$-point sampling in a finite simulation cell, the electron momentum states are quantized leading to errors in kinetic and exchange energies. (ii) Coulomb FS error, which is due to the issue that the long-range Coulomb interaction is artificially truncated by PBCs and therefore self-interaction and electrostatic image effects distort the interaction energy. (iii) Many-body correlation FS error mainly driven by the absence of long-wavelength collective excitations in small simulation cells. To correct FS errors, we used the Ewald summation with $\Gamma$-point and FS scaling extrapolation as shown in Fig. 3. The single-particle FS errors in the kinetic energy part of Hamiltonian can be corrected by twist averaging [61]. The almost perfect linear behavior of the VMC and DMC energies as a function of $1/N$ suggests that the single-particle FS errors are relatively small in our simulations.
Our VMC and DMC energies for double-layer FeSe at equilibrium configuration are obtained using two forms for two-body Jastrow functions. The pair correlation function $u_{F}(r)$ is more suitable in atomic chemical bond separation, while $u_{C}(r)$, which is large and positive for $r=0$ and goes to zero as $r\rightarrow\infty$, is useful when the atoms are at a distance greater than the variational parameter $b$. The difference between VMC energies at the thermodynamic limit obtained by the wave function $JSD$ and $u_{F}(r)$ and $u_{C}(r)$ is small. We propose the idea that using two different forms for the two-body Jastrow term and finding a good agreement between the VMC energies can be used as a test for the accuracy of wave function optimization. The DMC energies are almost independent of the form of Jastrow functions, as our results show (Fig. 3), since the nodal surface is not directly affected by the Jastrow function.
@@
Figure 3: (a) VMC and (b) DMC energy of double layer FeSe as a function of $1/N$, where $N$ is the number of electrons in the simulation cell, obtained using $J_{F}$ and $J_{C}$ two-body el-el Jastrow factors.
To investigate the effect of basis set size on our QMC energies, we increased the number of Fe-$d$ orbitals in the single-particle basis set of the Slater determinant and the Jastrow factor. The VMC and DMC energies of the double layer FeSe, which are calculated using the determinant and Jastrow base set of $8s6pnd1f,(n=4,6)$ and $3s3pnd,(n=1,3)$, respectively, for the iron atom, are listed in table 5. The energies are obtained for double-layer FeSe with Fe-Fe bond length of 3.69 Å  and interlayer separation of 5.83 Å. Increasing the number of Fe-$d$ orbitals in the Slater determinant basis set improves the ground-state energy more than in the Jastrow basis set. Compared with DMC, the VMC energies are more affected by the number of Fe-$d$ orbitals in the base set. More importantly, full optimization of the wave function reduces the effect of number of Fe-$d$ orbitals in the basis set as shown by the DMC-JSD energies (tab. 5). Comparison of the DMC-JSD energies of the smallest and largest basis sets shows that reaching the mHa/el accuracy is possible by full optimization of the wave function and without increasing the size of the basis set. The DMC-JSD energy difference between the smallest and the largest basis sets is less than 0.4 mH/el. Increasing the number of $d$ orbitals may improve the flexibility of the trial wave function and allows it to better represent subtle features in orbital hybridization and correlation, as this expansion led to better improvements in VMC energy, as the difference in VMC-JSD energy between the smallest and the largest basis sets is $\sim 2.8$ mHa/el.